silvernight269:

If measure angle 3 equals (4x+5y), angle 4 = (10x+10y), and angle 5 = (6x-20y), find the value of x, the value of y, and the measure of angle 8.

silvernight269:

1 attachment
SmokeyBrown:

So, from the picture, it looks like we have a set of interior opposite angles in angles 3 and 5. That means that if we add the measures of those angles together, we should get 180. Angles 3 and 4 are adjacent on the same line, so their measures will also add up to 180. So, we have Angle 3 + Angle 5 = Angle 3 + Angle 4 = 180 Now, let's bring the variables into this. 14x+15y=10x-15y 4x = -30y Now, we can substitute x for -30/4y, or -7.5y 10(-7.5y)-15y = 180 90y = 180 y = 2 And since 4x=30y, that means x=15. Now, you can just plug the values of x and y into the equations for each angle. Note that angle 4 is equal to angle 8, since they are corresponding angles. Or, we could use the fact that angle5 + angle8 = 180

silvernight269:

Ok

silvernight269:

i dont get it.

SmokeyBrown:

So we found that x is 15 and y is 2, right? And we have the equations for all of the angles, using x and y, right? So if we replace x with 15 and y with 2 in those equations, we'll find the measures of each of the angles

silvernight269:

Dude? can you explain what Smokey said for the first part?

dude:

Lol what part are you confused on? *highlight and repost it*

silvernight269:

when smokey replied after i posted the image

silvernight269:

@dude ?

dude:

Do you know what opposite interior angles are?

silvernight269:

yes

dude:

You know they add up to 180, right?

silvernight269:

yes

dude:

What did you not get? xD Do you know what adjacent means?

silvernight269:

yes. i dont get the equations. how did he get up to 14x+15y=10x-15y 4x = -30y?

dude:

Smokey added angle 3 (4x+5y) and angle 4 (10x+10y) to get 14x+15y Because they add up to 180 Angle 3 and angle 5 were added because they are adjacent (AKA supplementary; they both add up to 180) Since they both add up to 180, they should be the same (Angle 3+4)=(Angle 3+5) 14x + 15y = 10x-15y

silvernight269:

ok. then what.

silvernight269:

4x=-30y?

dude:

They subtracted 15y on both sides

silvernight269:

ok

silvernight269:

wait. cant i do 4x+30y=180?

dude:

No, you already set them equal to each other, setting that equal to 180 is not correct because 4x+30y is not either equation

silvernight269:

then i dont get it.

dude:

What 4x+30y is a modified version of it, they are not equal, you combined terms from 2 different equations, they can no longer be 180 What Smokey did was isolate x (by dividing by 4 on both sides) \(4x=-30y\) \(x=-\frac{30}{4}y=-7.5y\)

silvernight269:

ok

dude:

Then they substitute it into (angle 3+5) \(10(-7.5y)-15y = 180\)

silvernight269:

wait. so angle 3+5 is what?

dude:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dude Smokey added angle 3 (4x+5y) and angle 4 (10x+10y) to get 14x+15y Because they add up to 180 Angle 3 and angle 5 were added because they are adjacent (AKA supplementary; they both add up to 180) Since they both add up to 180, they should be the same (Angle 3+4)=(Angle 3+5) 14x + 15y = 10x-15y \(\color{#0cbb34}{\text{End of Quote}}\)

silvernight269:

ahhhhhhh. ok

silvernight269:

-15y=105